A ball is dropped from rest at the top of a 100 m tall building. At the same instant, a second ball is thrown upward from the base of the building with an initial velocity of 50 m/s. When will the two balls collide, and at what distance above the street? 


1
Expert's answer
2020-09-03T14:06:35-0400

The law of motion of the first ball

"y_1=y_0+v_{0y}t-\frac{gt^2}{2}\\
=100-\frac{9.8t^2}{2}=100-4.9t^2"

The law of motion of the second ball

"y_2=y_0+v_{0y}t-\frac{gt^2}{2}\\
=50t-\frac{9.8t^2}{2}=50t-4.9t^2"

At the instant when balls collide "y_1=y_2," so


"100-4.9t^2=50t-4.9t^2""t=2\:\rm s"

The distance above the street

"d=y_1(t)=y_2(t)=50\times 2-4.9\times 2^2=80.4\:\rm m"